# Universal bounds on optimization of free energy harvesting

### 20 Dec 2023 v3

@@ 1.4   Derivation and achievability of unconstrained optimization (6)
1.4.1   Construction of optimal control
@@
$R^{c}_{ji}(\kappa):=\kappa B_{ji}\qquad\qquad B_{ji}:=\frac{G_{ji}}{1+e^{\beta(f_{i}-f_{j}-{g}^{c}_{ji})}}=\kappa\frac{G_{ji}}{1+p_{j}^{*}/p_{i}^{*}}\,.$ (S21)
Here $\kappa\geq 0$ is an overall rate constant and $G$ is the matrix discussed in (S19). The particular choice of the topology encoded in $G$ is arbitrary, given our assumption that it is connected and contains all $n$ states which guarantees that $R^{c}$ is irreducible. It can be verified that the rate matrix $R^{c}(\kappa)$ defined in (S21) obeys LDB (2). It also satisfies detailed balance (DB) for the distribution $\bm{p}^{*}$,
$R^{c}_{ji}(\kappa){p}^{*}_{i}=R^{c}_{ij}(\kappa){p}^{*}_{j},$ (S22)
@@
We note that there are various other types of rate matrices that can be used in the construction, as long as they satisfy LDB and DB. However, the parameterization used in (S21) is common in the literature [52].
1.4.2   Achievability of the steady state $\bm{p}^{*}$
@@
$\frac{1}{\kappa}R^{b}\bm{\pi}({\kappa})=-\frac{1}{\kappa}R^{c}(\kappa)\bm{\pi}({\kappa})=\frac{1}{\kappa}R^{c}(\kappa)\left(\bm{p}^{*}-\bm{\pi}({\kappa})\right)=B(\bm{p}^{*}-\bm{\pi}({\kappa})).$ (S23)
@@
$\frac{1}{\kappa}B^{+}R^{b}\bm{\pi}({\kappa})=B^{+}B\left(\bm{p}^{*}-\bm{\pi}({\kappa})\right)\,.$ (S24)
Since $B$ is irreducible by construction, $B^{+}B=I-\bm{p}^{*}\boldsymbol{1}^{T}$. Plugging into (S24), we obtain
$\frac{1}{\kappa}B^{+}R^{b}\bm{\pi}({\kappa})=(I-\bm{p}^{*}\boldsymbol{1}^{T})(\bm{p}^{*}-\bm{\pi}({\kappa}))=\bm{p}^{*}-\bm{\pi}({\kappa}),$ (S25)
@@
$\left\|\bm{p}^{*}-\bm{\pi}({\kappa})\right\|=\frac{1}{\kappa}\|B^{+}R^{b}\bm{\pi}({\kappa})\|\leq\frac{1}{\kappa}\|B^{+}R^{b}\|\|\bm{\pi}({\kappa})\|\leq\frac{1}{\kappa}\|B^{+}R^{b}\|,$ (S26)
@@ 1.4.3   EPR vanishes
$\dot{\Sigma}(\kappa)=\frac{\kappa}{2}\sum_{i\neq j}(\pi_{j}(\kappa)B_{ij}-\pi_{i}(\kappa)B_{ji})\ln\frac{\pi_{j}(\kappa)B_{ij}}{\pi_{i}(\kappa)B_{ji}}\,.$ (S27)
Next, we use (S25) to write $\bm{p}^{*}=\bm{\pi}({\kappa})+\frac{1}{\kappa}B^{+}R^{b}\bm{\pi}({\kappa})$. Using the DB condition for $\bm{p}^{*}$ with respect to $R^{c}(\kappa)=\kappa B$ obtained in (S22), we rearrange terms to find:
$\pi_{i}(\kappa)B_{ji}-\pi_{j}(\kappa)B_{ij}=\frac{1}{\kappa}\left[B_{ij}\left(B^{+}R^{b}\bm{\pi}({\kappa})\right)_{j}-B_{ji}\left(B^{+}R^{b}\bm{\pi}({\kappa})\right)_{i}\right]$ (S28)
Plugging into (S27) gives
$\dot{\Sigma}(\kappa)=\frac{1}{2}\sum_{i\neq j}\left[B_{ji}\left(B^{+}R^{b}\bm{\pi}({\kappa})\right)_{i}-B_{ij}\left(B^{+}R^{b}\bm{\pi}({\kappa})\right)_{j}\right]\ln\frac{\pi_{j}(\kappa)B_{ij}}{\pi_{i}(\kappa)B_{ji}}.$ (S29)
@@
where we used $\lim_{\kappa\to\infty}\bm{\pi}({\kappa})=\bm{p}^{*}$ and the DB condition (S22) in the logarithmic factor.
Bacteriorhodopsin model
2.1   Details of model
@@
$\begin{gathered}g_{LK}=g_{KL}=g_{M_{1}L}=g_{LM_{1}}=g_{NM_{2}}=g_{M_{2}N}=g_{ON}=g_{NO}=g_{KO}=g_{OK}=0\\ g_{M_{2}M_{1}}=-g_{M_{1}M_{2}}=\Delta\mathrm{p}:=e\Delta\psi-(\ln 10)\beta^{-1}\Delta\text{pH}\qquad\text{joules}\end{gathered}$ (S33)
@@
$R_{ji}=\frac{\kappa_{ji}}{1+e^{-\Delta s^{\mathrm{tot}}_{ji}}}\,,$ (S35)
@@
$\Delta s^{\mathrm{tot}}_{ji}=\beta(f_{i}-f_{j}-g_{ji}+m_{ji})\,.$ (S36)
where $f_{i}$ refers to internal free energies in joules (S30), $\beta=1/k_{B}T$ and $k_{B}=1.38\times 10^{-23}\ \text{joules}\cdot\left({}^{\circ}\text{ K}\right)^{-1}$ is Boltzmann’s constant. As above $g_{ji}=-g_{ij}$ is the increase of free energy of the internal reservoir, while $m_{ji}=-m_{ij}$ is the decrease of free energy in the external source during the transition $i\to j$. In fact, $m_{ji}=0$ for all $i,j$ except the transition $O\leftrightarrow K$, which corresponds to photon absorption. The energy absorbed from the photon is $hc/\lambda$ joules, where $h$ is Planck’s constant, $c$ is speed of light, and $\lambda$ is the photon wavelength. We use a physiologically plausible value of 580 nm [53]. At this wavelength and temperature of $293^{\circ}$ K,
$m_{KO}=-m_{OK}=hc/\lambda\;\;\text{joules}\quad\approx\quad 84\,k_{B}T\,.$ (S37)
Observe that the transition $O\to K$ is highly irreversible ($r_{KO}\gg r_{OK}$). In fact, our results are essentially the same regardless of whether the transition rate for this step is computed using (S35) or made absolutely irreversible.
@@
$R=\,$ $\left[\begin{array}[]{rrrrrr}-2.40\times 10^{5}&1.70\times 10^{4}&0&0&0&1.28\times 10^{2}\\ 2.40\times 10^{5}&-2.57\times 10^{4}&1.68\times 10^{4}&0&0&0\\ 0&8.69\times 10^{3}&-1.99\times 10^{4}&2.35\times 10^{3}&0&0\\ 0&0&3.07\times 10^{3}&-2.77\times 10^{3}&3.01\times 10^{2}&0\\ 0&0&0&4.20\times 10^{2}&-4.75\times 10^{2}&3.37\times 10^{2}\\ 6.78\times 10^{-32}&0&0&0&1.74\times 10^{2}&-4.65\times 10^{2}\end{array}\right]\,.$ (S46)
@@
2.2   Details of numerical analysis in Fig. 1
@@
    Use the transition rates from (S35) to define the ‘total’ (baseline-and-control) rate matrix $R$ and numerically solve for $R\bm{\pi}=0$ to obtain $\bm{\pi}$, which is unique since $R$ is irreducible (all states are connected, see Fig. S3).
@@
    Use the transition rates from (S35) and values of $\bm{g}$ from (S33) to calculate the total harvesting rate as
    $\dot{\mathcal{G}}^{\mathrm{tot}}=\sum_{i,j}\pi_{i}R_{ji}g_{ji}\,.$ (S47)
@@
      Use the transition rates and free energy values of $\bm{f}$ from (S30) to fix the “baseline harvesting rate” and the “increase of system free energy” (Eq. (1)) functions:
      $\dot{\mathcal{G}}^{b}(\bm{p})$ $=\sum_{i,j}p_{i}R^{b}_{ji}g_{ji}$ (S48)
      ˙$\dot{\mathcal{F}}^{b}(\bm{p})$ $=\sum_{i,j}p_{i}R^{b}_{ji}(f_{j}+{\beta}^{-1}\ln p_{j}).$ (S49)
@@
2.3   Additional analysis: reprotonation step as control
@@
$\Lambda^{NO}:=\{(\bm{p},{\boldsymbol{J}}):\bm{p}\geq 0,\bm{1}^{T}\bm{p}=1,{\boldsymbol{J}}\geq 0,J_{ij}=0\text{ if }(i,j)\not\in\{{(N,O)},{(O,N)}\}\}\,.$ (S50)
@@
$R^{c}_{ON}=\frac{\kappa_{ON}}{1+e^{-\Delta s^{\mathrm{tot}}_{ON}}}\qquad\qquad R^{c}_{NO}=\frac{\kappa_{NO}}{1+e^{-\Delta s^{\mathrm{tot}}_{ON}}}$ (S51)
with all other $R^{c}_{ij}=0$. The baseline-and-control dynamics correspond to the actual bacteriorhodopsin system described in Sec. 2.1. See also Fig. S4 (left).
For Sections 2.3 and 2.4 in this SM, it is useful to define the current (i.e., net flux) as
$\mathcal{J}=J_{NO}^{*}-J_{ON}^{*}$ (S52)
@@
Figure S4: Depiction of choices for baseline and control for Sections 2.3 (left) and 2.4 (right) of the SM. In each case, the baseline (shaded blue) does not include the transition of reprotonation ($N\leftrightarrow O$) and proton-pumping ($M_{1}\leftrightarrow M_{2}$), respectively; instead, these are treated as control (shaded yellow). For both scenarios, as in the analysis done in the main text, note that no cyclic current circulates under baseline dynamics but does so under baseline-and-control (i.e. the actual) dynamics.
2.3.1   Constrained activity and affinity
@@
$\Lambda_{a}^{NO}:=\Lambda^{NO}\cap\{(\bm{p},{\boldsymbol{J}}):J_{NO}+J_{ON}\leq a\},$ (S53)
@@
Thus, constraining the activity as in (S53) limits the ability of the optimized system to amplify the current above the actual in vivo current (see top left plot in Fig. S5). In contrast, increasing $a$ approaches the optimal solution to the one shown for the $N\leftrightarrow O$ transition in the main text (Fig. 1), here shown in red. Note also that for strongly constrained activity, such as $a=10^{1}$, the optimal distribution around the in vivo value shifts in the opposite way with respect to the unconstrained case (blue against red bar-plots in the lower left of Fig. S5).
@@
$\Lambda_{\mathcal{A},a}^{NO}:=\Lambda_{a}^{NO}\cap\left\{(\bm{p},{\boldsymbol{J}}):J_{NO}e^{-\mathcal{A}}\leq J_{ON}\leq J_{NO}e^{\mathcal{A}}\right\}.$ (S54)
@@
$\left|\ln\frac{J_{ON}}{J_{NO}}\right|\leq\mathcal{A}\qquad\text{for some}\quad\mathcal{A}\geq 0\,.$ (S55)
@@
The combination of the constraint on dynamical activity and thermodynamic affinity is necessary to give physically meaningful results. To see why, imagine that only the thermodynamic affinity was constrained, as in (S55). Now consider some pair of a distribution $\bm{p}$ and flux vector ${\boldsymbol{J}}$ in our optimization problem (5), where ${\boldsymbol{J}}$ is restricted to have non-zero transitions only for $N\to O$ and $O\to N$, and satisfy the steady-state condition $\mathbb{B}{\boldsymbol{J}}=-R^{b}\bm{p}$. These fluxes can be increased as $J_{NO}\mapsto J_{NO}+\alpha,J_{ON}\mapsto J_{ON}+\alpha$ for $\alpha\geq 0$. This transformation doesn’t change the current across the transition $N\to O$ (S52), so the steady-state constraint is still valid for the same $\bm{p}$ (note that $\mathbb{B}{\boldsymbol{J}}$ only depends on the currents, i.e., net fluxes). Finally, by choosing $\alpha$ large enough, we can make the EPR term in (5) vanish and satisfy the affinity constraint (S55), since $\lim_{\alpha\to\infty}|\ln[(J_{NO}+\alpha)/(J_{ON}+\alpha)]|=0$. This shows that, lacking other constraints, the affinity constraint is vacuous, since it can be always satisfied by making the one-way fluxes sufficiently large. On the other hand, the combination of the activity and affinity constraints sets a bound on the fluxes.
2.3.2   Constrained kinetics

### 12 Mar 2024 v4

@@ I.4 Derivation and achievability of unconstrained optimization (11)
I.4.1 Construction of optimal control
@@
$R^{c}_{ji}(\kappa):=\kappa B_{ji}\qquad\qquad B_{ji}:=\frac{G_{ji}}{1+e^{\beta(f_{i}-f_{j}-{g}^{c}_{ji})}}=\kappa\frac{G_{ji}}{1+p_{j}^{*}/p_{i}^{*}}\,.$ (S1.21)
Here $\kappa\geq 0$ is an overall rate constant and $G$ is the matrix discussed in (S1.19). The particular choice of the topology encoded in $G$ is arbitrary, given our assumption that it is connected and contains all $n$ states which guarantees that $R^{c}$ is irreducible. It can be verified that the rate matrix $R^{c}(\kappa)$ defined in (S1.21) obeys LDB (6). It also satisfies detailed balance (DB) for the distribution $\bm{p}^{*}$,
$R^{c}_{ji}(\kappa){p}^{*}_{i}=R^{c}_{ij}(\kappa){p}^{*}_{j},$ (S1.22)
@@
We note that there are various other types of rate matrices that can be used in the construction, as long as they satisfy LDB and DB. However, the parameterization used in (S1.21) is common in the literature [76].
I.4.2 Achievability of the steady state $\bm{p}^{*}$
@@
$\frac{1}{\kappa}R^{b}\bm{\pi}({\kappa})=-\frac{1}{\kappa}R^{c}(\kappa)\bm{\pi}({\kappa})=\frac{1}{\kappa}R^{c}(\kappa)\left(\bm{p}^{*}-\bm{\pi}({\kappa})\right)=B(\bm{p}^{*}-\bm{\pi}({\kappa})).$ (S1.23)
@@
$\frac{1}{\kappa}B^{+}R^{b}\bm{\pi}({\kappa})=B^{+}B\left(\bm{p}^{*}-\bm{\pi}({\kappa})\right)\,.$ (S1.24)
Since $B$ is irreducible by construction, $B^{+}B=I-\bm{p}^{*}\boldsymbol{1}^{T}$. Plugging into (S1.24), we obtain
$\frac{1}{\kappa}B^{+}R^{b}\bm{\pi}({\kappa})=(I-\bm{p}^{*}\boldsymbol{1}^{T})(\bm{p}^{*}-\bm{\pi}({\kappa}))=\bm{p}^{*}-\bm{\pi}({\kappa}),$ (S1.25)
@@
$\left\|\bm{p}^{*}-\bm{\pi}({\kappa})\right\|=\frac{1}{\kappa}\|B^{+}R^{b}\bm{\pi}({\kappa})\|\leq\frac{1}{\kappa}\|B^{+}R^{b}\|\|\bm{\pi}({\kappa})\|\leq\frac{1}{\kappa}\|B^{+}R^{b}\|,$ (S1.26)
@@ I.4.3 EPR vanishes
$\dot{\Sigma}(\kappa)=\frac{\kappa}{2}\sum_{i\neq j}(\pi_{j}(\kappa)B_{ij}-\pi_{i}(\kappa)B_{ji})\ln\frac{\pi_{j}(\kappa)B_{ij}}{\pi_{i}(\kappa)B_{ji}}\,.$ (S1.27)
Next, we use (S1.25) to write $\bm{p}^{*}=\bm{\pi}({\kappa})+\frac{1}{\kappa}B^{+}R^{b}\bm{\pi}({\kappa})$. Using the DB condition for $\bm{p}^{*}$ with respect to $R^{c}(\kappa)=\kappa B$ obtained in (S1.22), we rearrange terms to find:
$\pi_{i}(\kappa)B_{ji}-\pi_{j}(\kappa)B_{ij}=\frac{1}{\kappa}\left[B_{ij}\left(B^{+}R^{b}\bm{\pi}({\kappa})\right)_{j}-B_{ji}\left(B^{+}R^{b}\bm{\pi}({\kappa})\right)_{i}\right]$ (S1.28)
Plugging into (S1.27) gives
$\dot{\Sigma}(\kappa)=\frac{1}{2}\sum_{i\neq j}\left[B_{ji}\left(B^{+}R^{b}\bm{\pi}({\kappa})\right)_{i}-B_{ij}\left(B^{+}R^{b}\bm{\pi}({\kappa})\right)_{j}\right]\ln\frac{\pi_{j}(\kappa)B_{ij}}{\pi_{i}(\kappa)B_{ji}}.$ (S1.29)
@@
where we used $\lim_{\kappa\to\infty}\bm{\pi}({\kappa})=\bm{p}^{*}$ and the DB condition (S1.22) in the logarithmic factor.
II Bacteriorhodopsin model
II.1 Details of model
@@
$\begin{gathered}g_{LK}=g_{KL}=g_{M_{1}L}=g_{LM_{1}}=g_{NM_{2}}=g_{M_{2}N}=g_{ON}=g_{NO}=g_{KO}=g_{OK}=0\\ g_{M_{2}M_{1}}=-g_{M_{1}M_{2}}=\Delta\mathrm{p}:=e\Delta\psi-(\ln 10)\beta^{-1}\Delta\text{pH}\qquad\text{joules}\end{gathered}$ (S2.33)
@@
$R_{ji}=\frac{\kappa_{ji}}{1+e^{-\Delta s^{\mathrm{tot}}_{ji}}}\,,$ (S2.35)
@@
$\Delta s^{\mathrm{tot}}_{ji}=\beta(f_{i}-f_{j}-g_{ji}+m_{ji})\,.$ (S2.36)
where $f_{i}$ refers to internal free energies in joules (S2.30), $\beta=1/k_{B}T$ and $k_{B}=1.38\times 10^{-23}\ \text{joules}\cdot\left({}^{\circ}\text{ K}\right)^{-1}$ is Boltzmann’s constant. As above $g_{ji}=-g_{ij}$ is the increase of free energy of the internal reservoir, while $m_{ji}=-m_{ij}$ is the decrease of free energy in the external source during the transition $i\to j$. In fact, $m_{ji}=0$ for all $i,j$ except the transition $O\leftrightarrow K$, which corresponds to photon absorption. The energy absorbed from the photon is $hc/\lambda$ joules, where $h$ is Planck’s constant, $c$ is speed of light, and $\lambda$ is the photon wavelength. We use a physiologically plausible value of 580 nm [77]. At this wavelength and temperature of $293^{\circ}$ K,
$m_{KO}=-m_{OK}=hc/\lambda\;\;\text{joules}\quad\approx\quad 84\,k_{B}T\,.$ (S2.37)
Observe that the transition $O\to K$ is highly irreversible ($r_{KO}\gg r_{OK}$). In fact, our results are essentially the same regardless of whether the transition rate for this step is computed using (S2.35) or made absolutely irreversible.
@@
$R=\,$ $\left[\begin{array}[]{rrrrrr}-2.40\times 10^{5}&1.70\times 10^{4}&0&0&0&1.28\times 10^{2}\\ 2.40\times 10^{5}&-2.57\times 10^{4}&1.68\times 10^{4}&0&0&0\\ 0&8.69\times 10^{3}&-1.99\times 10^{4}&2.35\times 10^{3}&0&0\\ 0&0&3.07\times 10^{3}&-2.77\times 10^{3}&3.01\times 10^{2}&0\\ 0&0&0&4.20\times 10^{2}&-4.75\times 10^{2}&3.37\times 10^{2}\\ 6.78\times 10^{-32}&0&0&0&1.74\times 10^{2}&-4.65\times 10^{2}\end{array}\right]\,.$ (S2.46)
@@
II.2 Details of numerical analysis in Fig. 2
@@
    Use the transition rates from (S2.35) to define the ‘total’ (baseline-and-control) rate matrix $R$ and numerically solve for $R\bm{\pi}=0$ to obtain $\bm{\pi}$, which is unique since $R$ is irreducible (all states are connected, see Fig. S7).
@@
    Use the transition rates from (S2.35) and values of $\bm{g}$ from (S2.33) to calculate the total harvesting rate as
    $\dot{\mathcal{G}}^{\mathrm{tot}}=\sum_{i,j}\pi_{i}R_{ji}g_{ji}\,.$ (S2.47)
@@
      Use the transition rates and free energy values of $\bm{f}$ from (S2.30) to fix the “baseline harvesting rate” and the “increase of system free energy” (Eq. (2)) functions:
      $\dot{\mathcal{G}}^{b}(\bm{p})$ $=\sum_{i,j}p_{i}R^{b}_{ji}g_{ji}$ (S2.48)
      ˙$\dot{\mathcal{F}}^{b}(\bm{p})$ $=\sum_{i,j}p_{i}R^{b}_{ji}(f_{j}+{\beta}^{-1}\ln p_{j}).$ (S2.49)
@@
II.3 Additional analysis: reprotonation step as control
@@
$\Lambda^{NO}:=\{(\bm{p},{\boldsymbol{J}}):\bm{p}\geq 0,\bm{1}^{T}\bm{p}=1,{\boldsymbol{J}}\geq 0,J_{ij}=0\text{ if }(i,j)\not\in\{{(N,O)},{(O,N)}\}\}\,.$ (S2.50)
@@
$R^{c}_{ON}=\frac{\kappa_{ON}}{1+e^{-\Delta s^{\mathrm{tot}}_{ON}}}\qquad\qquad R^{c}_{NO}=\frac{\kappa_{NO}}{1+e^{-\Delta s^{\mathrm{tot}}_{ON}}}$ (S2.51)
with all other $R^{c}_{ij}=0$. The baseline-and-control dynamics correspond to the actual bacteriorhodopsin system described in Sec. II.1. See also Fig. S8 (left).
For Sections II.3 and II.4 in this SM, it is useful to define the current (i.e., net flux) as
$\mathcal{J}=J_{NO}^{*}-J_{ON}^{*}$ (S2.52)
@@
Figure S8: Depiction of choices for baseline and control for Sections II.3 (left) and II.4 (right) of the SM. In each case, the baseline (shaded blue) does not include the transition of reprotonation ($N\leftrightarrow O$) and proton-pumping ($M_{1}\leftrightarrow M_{2}$), respectively; instead, these are treated as control (shaded yellow). For both scenarios, as in the analysis done in the main text, note that no cyclic current circulates under baseline dynamics but does so under baseline-and-control (i.e. the actual) dynamics.
II.3.1 Constrained activity and affinity
@@
$\Lambda_{a}^{NO}:=\Lambda^{NO}\cap\{(\bm{p},{\boldsymbol{J}}):J_{NO}+J_{ON}\leq a\},$ (S2.53)
@@
Thus, constraining the activity as in (S2.53) limits the ability of the optimized system to amplify the current above the actual in vivo current (see top left plot in Fig. S9). In contrast, increasing $a$ approaches the optimal solution to the one shown for the $N\leftrightarrow O$ transition in the main text (Fig. 2), here shown in red. Note also that for strongly constrained activity, such as $a=10^{1}$, the optimal distribution around the in vivo value shifts in the opposite way with respect to the unconstrained case (blue against red bar-plots in the lower left of Fig. S9).
@@
$\Lambda_{\mathcal{A},a}^{NO}:=\Lambda_{a}^{NO}\cap\left\{(\bm{p},{\boldsymbol{J}}):J_{NO}e^{-\mathcal{A}}\leq J_{ON}\leq J_{NO}e^{\mathcal{A}}\right\}.$ (S2.54)
@@
$\left|\ln\frac{J_{ON}}{J_{NO}}\right|\leq\mathcal{A}\qquad\text{for some}\quad\mathcal{A}\geq 0\,.$ (S2.55)
@@
The combination of the constraint on dynamical activity and thermodynamic affinity is necessary to give physically meaningful results. To see why, imagine that only the thermodynamic affinity was constrained, as in (S2.55). Now consider some pair of a distribution $\bm{p}$ and flux vector ${\boldsymbol{J}}$ in our optimization problem (10), where ${\boldsymbol{J}}$ is restricted to have non-zero transitions only for $N\to O$ and $O\to N$, and satisfy the steady-state condition $\mathbb{B}{\boldsymbol{J}}=-R^{b}\bm{p}$. These fluxes can be increased as $J_{NO}\mapsto J_{NO}+\alpha,J_{ON}\mapsto J_{ON}+\alpha$ for $\alpha\geq 0$. This transformation doesn’t change the current across the transition $N\to O$ (S2.52), so the steady-state constraint is still valid for the same $\bm{p}$ (note that $\mathbb{B}{\boldsymbol{J}}$ only depends on the currents, i.e., net fluxes). Finally, by choosing $\alpha$ large enough, we can make the EPR term in (10) vanish and satisfy the affinity constraint (S2.55), since $\lim_{\alpha\to\infty}|\ln[(J_{NO}+\alpha)/(J_{ON}+\alpha)]|=0$. This shows that, lacking other constraints, the affinity constraint is vacuous, since it can be always satisfied by making the one-way fluxes sufficiently large. On the other hand, the combination of the activity and affinity constraints sets a bound on the fluxes.
II.3.2 Constrained kinetics